a rectangular parking lot has a length that is 8 yards greater than the width. The area of the parking lot is 180 square yards. Find the length and the width.

Answer by checkley79(3341)

L*W=AREA
L=W+8
(W+8)W=180
W^2+8W=180
W^2+8W-180=0
(W-10)(W+18)=0
W-10=0
W=10 ANS. FOR THE WIDTH.
L=10+8=18 ANS. FOR THE LENGTH.
PROOF:
18*10=180
180=180